OpenStudy (anonymous):

why this summation ∑j=1to n+1((n+1)!/(n+1−j)!(j)!)pj(1−p)n+1−j = (1-(1-p) n+1 )

OpenStudy (anonymous):

\[\sum_{j=0}^{n+1}((n+1)!/(n+1−j)!(j)!)p ^{j} (1−p)^{ n+1−j }\] =(1-(1-p) \[^{n+1}\] )

OpenStudy (anonymous):

=1/(1-p)^n??????

OpenStudy (anonymous):

=1/(1-p)^(n+1) ???????

OpenStudy (anonymous):

1-(1-p)^(n+1)

OpenStudy (anonymous):